Find both the x-intercept and the y-intercept of the line given by the equation.
8x+4y+5=0

Found 2 solutions by nerdybill, solver91311:
Answer by nerdybill(7384)

8x+4y+5=0
.
To find the y-intercept, put it into the "slope-intercept" form:
y = mx + b
.
8x+4y+5=0
8x+4y= -5
4y= -8x - 5
y = -2x - 5/4
.
y-intercept @ (0, -5/4)
.
To find the x-intercept, set y to zero and solve for x:
8x+4y+5=0
8x+4(0)+5=0
8x+5=0
8x = -5
x = -5/8
.
x-intercept @ (-5/8, 0)

Answer by solver91311(24713)

The x-intercept is at the point (a, 0) where a is the value of x when 0 is substituted for y. Conversely, the y-intercept is at the point (0, b) where b is the value of y when 0 is substituted for x.